"ACS Transition to Operate1 - Ops Request 17802-0 was completed at 173/21:11:12, successfully transitioning ACS from Suspend to its Operate1 state. In this state, ACS normal engineering data collection can be observed."
"On Monday, 19 June 2006, at 1:15 pm EDT (17:15 UT), the Advanced Camera for Surveys (ACS) issued status buffer messages indicating that the +15V and +5V power supply voltages in the CCD Electronics Box (CEB) were above their high limits, causing the ACS to suspend. This event occurred in a period with no ACS commanding and outside the SAA. A dump of the relevant data showed that a total of 36 CEB items exceeded limits at the time of the event."
The ACS suspended at 170/17:15:25z. An Ops Briefing was held at 6 pm on June 19, 2006. At 170/17:15:25 the ACS 715 and ACS 707 status buffer (STB) messages were received indicating the ACS WFC CEB analog signal processing +15 and + 5 volt power supply voltages were out of limits high which resulted in the ACS suspend. Detailed analysis of the event is underway with a tiger team meeting planned for Tuesday.
Flash Report: Status Briefing - ACS Suspend Status Update
A follow-up briefing was held at 1 pm on June 20, 2006 to review the status of the analysis on the ACS suspend event.
The preliminary data analysis of the event was presented as well as a review of the data collection process and the A/D FIFO Error handling. The forward plan for the STScI regarding the present SMS and near-term SMSs was discussed. It was agreed to intercept the science timeline with an SMS without ACS effective 8am EDT Thursday and to develop next week's SMS without ACS activities. The STScI will be ready to provide an intercept SMS with ACS observations when needed. The options for on-orbit tests to troubleshoot side 1 were also discussed.
It was agreed to reconvene on Thursday at 1pm to review the hardware assessment and the options for on-orbit activities.